OpenStudy (anonymous):

A 1.94-kg object moving to the right at 5.90 m/s collides head-on with a 3.89-kg object that is initially at rest. After the collision, the 1.94-kg object is moving to the left at 1.21 m/s. Calculate the velocity of the 3.89-kg object after the collision. Calculate the energy lost in the collision. What is the coefficient of restitution for these objects?

OpenStudy (anonymous):

well momentum = mv m = mass v = velocity momentum is conserved so the momentum before collision of the two particles is equal to the momentum after collision initial momentum of the 1.94 kg object is: mv = (1.94)(5.9) = 11.4 initial momentum of the 3.89 kg object is 0 because it is at rest so the total initial momentum is 11.4+0=11.4 final momentum on the 1.94 kg object is: mv = (1.94)(-1.21) = -2.35 final momentum of the 3.89 kg object is: mv = 3.89v so the total final momentum should be 3.89v-2.35=11.4 so solving for v gives: v = 3.53 ms^-1 coefficient of restitution: \[(3.53-1.21)\div(5.9-0)\] = 0.39